OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 4, an = an-1 + 8 OPTIONS ARE 0, 8, 16, 24, 32, 40 12, 20, 28, 36, 44, 52 4, 12, 20, 28, 36, 44 4, 8, 16, 24, 32, 40 PLEASE HELP I GIVE MEDALS

OpenStudy (raden):

just put n=2,3,4,5,6

OpenStudy (anonymous):

@RadEn where

OpenStudy (raden):

im here :)

OpenStudy (raden):

have u got it ?

OpenStudy (raden):

well, i'll give example ... a_1 = 4 (known) a_2 = a_(2-1) + 8 = a_1 + 8 = 4+8 = 12 a_3 = a_(3-1) + 8 = a_2 + 8 = 12+8 = 20 so on ....

OpenStudy (anonymous):

eek

OpenStudy (raden):

looks it is an arithmetic sequence with difference is 8 just add by 8 to the next number

OpenStudy (anonymous):

28 36 44 ?

OpenStudy (raden):

yep

OpenStudy (anonymous):

@RadEn what about a1 = 6, an = 2 • an-1

OpenStudy (raden):

do same idea, just put n=2 to get a_2 so, a_2 = 2a_(2-1) = 2a_1 = 2*6 = 12 and so on ...

OpenStudy (raden):

looks like a geometri sequence with the ratio = 2

OpenStudy (anonymous):

72

OpenStudy (anonymous):

what do you keep multiplying ?

OpenStudy (anonymous):

and for the last one does it start with 4,then 12 or is it starting with 12

OpenStudy (raden):

just multiply by 2 to the next term, they are be 6,12,24,48,96,192,...

OpenStudy (raden):

i mean to get the next term

OpenStudy (anonymous):

so for the last one it was 4 , 12 ,20, 28...

OpenStudy (raden):

6,12,24,48,96,192 the first six